# Main-belt and Trojan Asteroid Phase Curves from the ATLAS Survey

### v1

@@ Acknowledgements
The databases constructed in this work will be available in their entirety via the Queen’s University Belfast PURE open data repository (https://pure.qub.ac.uk/en/datasets/), and will also be submitted to the NASA PDS Small Bodies Node (https://pds-smallbodies.astro.umd.edu). They may also be obtained via reasonable e-mail request to the lead authors.
@@ Appendix B Comparing $HG$ and $H_{12}G_{12}$
Figure 17: Plots of B89 model $HG$ phase curves (blue curves), which were generated with a range of $0\leq G\leq 1$. Each $HG$ phase curve was then fitted with the P16 $H_{12}G_{12}$ model (orange curves). The two models have best agreement at large phase angles, however, they vary strongly at lower phase angles leading to differences in absolute magnitude.
@@ Appendix C Jupiter Trojan KS bootstrap tests
Figure 18: Distributions of $D$ values for KS tests comparing $H$, $G$ and $H_{c}-H_{o}$ colour of the L4 and L5 Jupiter Trojans in the SAP dataset (left, middle and right panels respectively; compare to figure 15 CDFs). The vertical blue line indicates the median $D$ of each distribution and the red line indicates the critical value $D_{\text{crit}}$ for rejection of the null hypothesis that the two distributions are the same (equation 10).

### v2

@@ Acknowledgements
The data bases constructed in this work are available in their entirety online as supplementary data or via the Queen’s University Belfast PURE open data repository (https://doi.org/10.17034/db29dc08-5e21-403f-93f3-d60073b97916). They may also be obtained via reasonable e-mail request to the lead authors.
Supporting Information
Supplementary data are available at MNRAS online.
Table 1. Overview of the observations made by ATLAS and collected by rockAtlas, for each asteroid considered in this study.
Table 2. Results of the phase curve fitting algorithm to determine phase curves for each apparition of the objects from Table 1.
Table 3. The final B89 phase curve parameters for the selected objects in Tables 1, as determined from the individual apparitions in Table 2.
Table 4. The final $H_{12}G_{12}$ P16 phase curve parameters for the selected objects from Table 1, as determined from the individual apparitions in Table 2.
Please note: Oxford University Press is not responsible for the content or functionality of any supporting materials supplied by the authors. Any queries (other than missing material) should be directed to the corresponding author for the article.
@@ Appendix B Comparing $HG$ and $H_{12}G_{12}$
Figure 17: Plots of B89 model $HG$ phase curves, which were generated with a range of $0\leq G\leq 1$. Each $HG$ phase curve was then fitted with the P16 $H_{12}G_{12}$ model. The two models have best agreement at large phase angles, however, they vary strongly at lower phase angles leading to differences in absolute magnitude.
@@ Appendix C Jupiter Trojan KS bootstrap tests
Figure 18: Distributions of $D$ values for KS tests comparing $H$, $G$ and $H_{c}-H_{o}$ colour of the L4 and L5 Jupiter Trojans in the SAP dataset (left, middle and right panels respectively; compare to figure 15 CDFs). The vertical dashed line indicates the median $D$ of each distribution and the vertical solid line indicates the critical value $D_{\text{crit}}$ for rejection of the null hypothesis that the two distributions are the same (equation 10).